OpenStudy (anonymous):

Decide whether the set of all real numbers of the form m+nroot(2) where m is in Z and n is E (E is the ring of all even numbers) is a integral domain (or a field) with respect to the usual operations of addition and multiplication. If it is not an integral domain (or a field), state at least one condition that fails to hold.

OpenStudy (amistre64):

Z + 2n sqrt(2) hmmm, what are the conditions that it needs to satisfy?

OpenStudy (amistre64):

additive associativity and commutativity dont seem to be an issue 0+0.rt2 is fine for an additive identity z + 2n.rt2 -z-2n.rt2 ; additive inverse works ---------- 0 + 0.rt2

OpenStudy (amistre64):

(1+0.rt2) would work as a mult identity

OpenStudy (amistre64):

(a+b.rt2)(x+y.rt2) = ax + 4by + (ay+bx)rt2 is there an ax+by, that is not even? ax + 2(2kn) .... if ax is odd we might have a fail on this since its not closed

OpenStudy (anonymous):

Yes those conditions. Also Identity element

OpenStudy (amistre64):

(a+b.rt2)(x+y.rt2) = ax + 4by + (ay+bx)rt2 is there an ay+bx, that is not even? b and y are always even, so yeah, thats fine

OpenStudy (amistre64):

do we have an inverse for multiplication?

OpenStudy (amistre64):

ax + 4by + (ay+bx)rt2 = 1 + 0.rt2

OpenStudy (amistre64):

ay + bx = 0; ay = -bx is fine since even numbers have inverses ax + 4by = 1, hmmm

OpenStudy (amistre64):

4by is even, is there an ax that is odd?

OpenStudy (amistre64):

or at least with the conbditions that ay = -bx that is

OpenStudy (amistre64):

0s tend to cause troubles, so we should prolly work those in

OpenStudy (anonymous):

Okay Im reading them.

OpenStudy (amistre64):

(a+0.rt)(x+0.rt) = 1 ax + 0 + (0+0)rt2 = 1 we would need a fraction, wouldnt we

OpenStudy (amistre64):

a in Z x in Z ax = 1 is not possible

OpenStudy (anonymous):

yes, so basically it is not an integer because it will not equal 1.

OpenStudy (amistre64):

(3+0.rt2)(x+0.rt2) = (1+0.rt2) has no \(x\in Z\) solution; x=1/3

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

that does not state which element does not hold.

OpenStudy (amistre64):

it shows that there is no inverse for at least 1 element of the group. the element: (3+0.rt2) is of the form m=3 is an integer, n=0 is an even number the multiplicative identity is the element: (1+0.rt) (a+b.rt)(1+0.rt) = (a+b.rt)(1) = (a+b.rt) therefore the inverse of multiplcation would have to satisfy (a+b.rt) (x+y.rt) = (1+0.rt2) for every element of the stated structure since: (3+0.rt2) (x+y.rt2) = (1 + 0.rt2) cannot be satisfied ... there is at least one element that has no inverse in multiplication

OpenStudy (amistre64):

(3+0.rt2) (x+y.rt2) = (1 + 0.rt2) (3x + 3y.rt2) = (1 + 0.rt2) when 3x = 1, and when 3y = 0 since y=0 is fine and dandy, we would need to find an integer x, such that 3x = 1

OpenStudy (amistre64):

if we leave 3 just a random integer; then for any a,x in Z; ax = 1 would have to be satisfied

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is this a ring?

OpenStudy (amistre64):

a ring stops at distrbutive property and ignores the mult id and inv right?

OpenStudy (amistre64):

a ring is: +retrice +id +inv *distr over +

OpenStudy (amistre64):

+assoc...

OpenStudy (amistre64):

some texts say: *assoc..

OpenStudy (amistre64):

either way, it only fails for a *inv, so i would say it satisfies a ring